# Large-Scale Text-to-Image Model with Inpainting is a Zero-Shot Subject-Driven Image Generator

### v1

@@ 1 Introduction
  • To the best of our knowledge, we are the first to approach zero-shot subject-driven text-to-image generation as an inpainting task, offering a new perspective by highlighting the diptych generation capabilities of FLUX.
@@ 2.3 Subject-Driven Image Generation
The second group [50, 22, 47, 30, 46, 26, 31] addresses the limitations of the first by introducing image prompting, a framework that incorporates a reference image to guide the generated output. In this approach, a specialized image enco der [34] is utilized to support an additional image input alongside the text prompt, allowing the diffusion model to capture the characteristics of the reference subject while adhering to the context described by the text prompt. This approach enables zero-shot subject-driven text-to-image generation; however, it often lacks fidelity in the synthesized subject. We propose a novel inpainting-based zero-shot approach for subject-driven text-to-image generation using a large-scale TTI model with an inpainting module.
@@ 4.2 Baseline Comparisons
We compare our method to previous zero-shot subject-driven text-to-image methods with encoder-based image prompting, including ELITE [47], BLIP-Diffusion [22], Kosmos-G [30], Subject-Diffusion [26], IP-Adapter [50], MS-Diffusion [46], and $\lambda$-Eclipse [31]. The details of these models are provided in Appendix A.
@@
Human Preference Study We confirm the outstanding performance of our method in terms of human perception through a human preference study. We conduct a paired comparisons of our method with each baseline from two perspectives: subject alignment and text alignment. Using Amazon Mechanical Turk, we collected $450$ responses from $150$ participants for each baseline and each perspective. As shown in Tab. 1, Diptych Prompting outperforms all baselines by a large margin ($p<0.01$ in the Wilcoxon signed-rank test), which is consistent with the qualitative results. Detailed information and full instructions about our human preference study are included in Appendix C.
@@
IP-Adapter [50$\ddagger$ FLUX 0.561 0.725 0.351
Diptych Prompting FLUX 0.688 0.758 0.345
@@ 4.3 Ablation Studies
We also assess the impact of varying the rescaling factor $\lambda$ on reference attention enhancement. Rescaling attention weights between the right panel query and the left panel key helps to capture fine details, thereby improving subject alignment. However, using too high values introduces excessive inductive bias, causing abnormal attention weights that negatively impact performance. Qualitative transitions with respect to $\lambda$ can be verified in Appendix F.
@@ 4.4 Applications
Stylized Image Generation We extend our method beyond subject images and perform stylized image generation using various style images as references. Using style images and prompts in StyleDrop [41], we employ Diptych Prompting as same, but replace the subject name with the term ‘style’ in diptych text and without attention enhancement ($\lambda=1$) for referencing only the stylistic elements except the content. Diptych Prompting successfully generates the stylistic image reflecting the style of the reference as shown in Fig. 6, and the quantitative comparisons are available in Appendix G.
@@ Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. 2022R1A3B1077720], Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University)], and the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024.
Appendix
Appendix A Baselines
@@
Figure A1: DreamBooth Comparisons. Quantitative comparisons to DreamBooth-LoRA with various rank values.
Appendix B Subject-Driven Text-to-Image Generation
@@ B.2 Comparison with Fine-Tuning-Based Method
To provide a more comprehensive comparison, we also compare with DreamBooth [39], a representative fine-tuning-based method. For efficient training, we attach a LoRA adapter to the pre-trained FLUX and perform fine-tuning by training only the LoRA adapter while freezing the FLUX. We train for 300 steps using the Adam optimizer with a learning rate of $1\times 10^{-4}$. Additionally, to compare different fine-tuning model capacities, we adjusted the rank of the LoRA adapter and conducted comparative experiments using the same metrics (DINO, CLIP-I, CLIP-T). The results are presented in Fig. A1, where our Diptych Prompting demonstrates superior performance across various model capacities.
@@ B.3 Additional Results
We include additional samples of Diptych Prompting in Fig. A2 and Fig. A3 for diverse objects and contexts. As demonstrated in the results, our methodology achieves high-quality image generation and satisfies both subject alignment and text alignment in a zero-shot manner by leveraging FLUX’s capabilities. Notably, this is accomplished without any specialized training for subject-driven text-to-image generation. We also note that the fine details in the target subject are well reflected in the generated results, even for challenging subjects that previous zero-shot methods struggled with (e.g., robot toy, ‘Bon appetit’ bowl).
Appendix C Human Preference Study
@@ Appendix D Diptych Generation
Table A1: Diptych Generation Comparisons. Quantitative comparisons of the diptych generation capabilities of various TTI models based on the total number of parameters, including the autoencoder, main network, and text encoder.
@@
Our framework relies on the emerging property of the large-scale TTI model, FLUX, particularly its strong understanding of diptych property and the ability to represent diptych accurately. We verify this by synthesizing a total of $2100$ diptychs, using $20$ objects, each with a pair of two random prompts for each panel among $15$ prompts, and comparing the diptych generation performance with those of other previous TTI models. The prompt for diptych generation follows the setup mentioned in Sec. 3.1 of the main paper. We assessed the quality of each diptych by evaluating the interrelation and text alignment of each panel. This is measured through splitting the generated image in half and measuring DINO and CLIP-I scores between each panel, as well as the CLIP-T score between each panel and its description. The results are shown in Tab. A1, in which the diptych generation performance and total number of parameters including the autoencoder, main network, and text encoders are reported. These results exhibit the superior diptych generation capability of FLUX, where smaller models are insufficient. This allows us to extend to inpainting and propose a zero-shot subject-driven text-to-image generation method via diptych inpainting-based interpretation.
Appendix E Background Removal Ablation
@@
We provide additional samples for the ablation study conducted with and without the background removal process $G_{\text{seg}}$ in Fig. A4. Consistent with the findings in the main paper, including the background leads to content leakage, where irrelevant elements such as background, pose, and location are mirrored in the generated results. This hinders the accurate reflection of the desired context described by the text and reduces diversity in pose and location. In contrast, removing the background and retaining only the subject information in the reference image on the left panel allows the generated outputs to better align with the desired context while exhibiting greater diversity in pose and location.”
Appendix F Reference Attention Enhancement Ablation
@@
We further present the actual sample quality variations according to the reference attention rescaling factor $\lambda$ values to support the quantitative ablations in the main paper. These variations are visualized in Fig. A5. As seen in the qualitative results, the absence of reference attention enhancement ($\lambda=1.0$) can lead to a loss of fine details of the subject, resulting in subtle discrepancies such as the left eye of the backpack dog, the patch on its right eye, the fur color on the dog’s face, or the texture of the bear plushie’s fur. As the $\lambda$ value increases, these missed details are better preserved, leading to improved subject alignment performance. However, excessive enhancement can negatively impact the quality of the generated images, causing the subject to appear slightly blurred or exhibit minor color shifts.
Appendix G Stylized Image Generation
@@
For stylized image generation, Diptych Prompting places the style image in the left panel and inpaints the right panel using the text prompt “A diptych with two side-by-side images of same style. On the left, {original image description}. On the right, replicate this style exactly but as {target image description} without attention enhancement ($\lambda=1.0$) for referencing only the stylistic elements except the content. Additional samples are provided in Fig. A6. Beyond the qualitative results, we also include quantitative comparisons using the same metrics (DINO, CLIP-T, CLIP-I) applied to a total of $2000$ generated images in Tab. A2. These images include $4$ samples per prompt and per style image, across $25$ prompts and $20$ style images collected from previous work [41]. As shown in the result, our method demonstrates comparable results to existing zero-shot style transfer methods specialized in stylized image generation, further proving the versatility of our approach.
@@
Table A2: Stylized Image Generation Comparisons. Quantitative comparisons of stylized image generation with previous zero-shot methods.
Appendix H Subject-Driven Image Editing
Diptych Prompting is extended to the subject-driven image editing by placing the reference subject image on the left panel and the editing target image on the right panel in the incomplete diptych. By masking only the desired area in the right panel and applying diptych inpainting, the reference subject from the left panel is generated in the masked region on the right panel, resulting in the subject-driven image editing. Following the previous work [49], we conduct the subject-driven image editing with selected images from a subset of the MSCOCO [24] validation dataset, in which each image contains a bounding box and the bounding box is smaller than half of image size. We applied masking to the inside of the bounding box, enabling the generation of the reference subject within the specified region. More samples of various subjects and editing target images are available in Fig. A7.
Appendix I Limitations
@@
Figure A2: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting.
Figure A3: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting..
Figure A4: $\bm{G_{\text{seg}}}$ Ablation. Qualitative comparisons with and without the background removal process.
Figure A5: $\bm{\lambda}$ Ablation. Qualitative transitions according to the varying $\lambda$ values. we control the $\lambda$ from $1.0$ (without reference attention enhancement) to $1.5$. For a detailed view, please zoom in.
Figure A6: Stylized Image Generation. More samples of stylized image generation using Diptych Prompting.
Figure A7: Subject-Driven Image Editing. More samples of subject-driven image editing using Diptych Prompting.

### v2

@@ 1 Introduction
  • We propose a novel inpainting-based zero-shot subject-driven text-to-image generation approach without further training, offering a new perspective by highlighting the inherent diptych generation capabilities of FLUX.
@@ 2.3 Subject-Driven Image Generation
The second group [55, 24, 51, 32, 50, 28, 33] addresses these limitations through image prompting, which utilizes a specialized image encoder [36] to incorporate a reference image alongside the text prompt to guide the generated output. Such methods enable zero-shot manner; yet, it often lacks target subject alignment. Other notable approaches [56, 17, 52] fine-tune the TTI model for joint-set image generation [2, 49, 46], and extend this to subject-driven generation or editing through inpainting. However, they still face training constraints, such as costs for dataset construction and training. Leveraging the inherent capability of a recent TTI model with an inpainting module, we propose a novel zero-shot inpainting-based approach without additional training.
@@ 4.2 Baseline Comparisons
We compare our method to previous zero-shot subject-driven text-to-image methods with encoder-based image prompting, including ELITE [51], BLIP-Diffusion [24], Kosmos-G [32], Subject-Diffusion [28], IP-Adapter [55], MS-Diffusion [50], and $\lambda$-Eclipse [33]. The details of these models are provided in appendix.
@@
Human Preference Study We confirm the outstanding performance of our method in terms of human perception through a human preference study. We conduct a paired comparisons of our method with each baseline from two perspectives: subject alignment and text alignment. Using Amazon Mechanical Turk, we collected $450$ responses from $150$ participants for each baseline and each perspective. As shown in Tab. 1, Diptych Prompting outperforms all baselines by a large margin ($p<0.01$ in the Wilcoxon signed-rank test), which is consistent with the qualitative results. Detailed information and full instructions about our human preference study are included in appendix.
@@
IP-Adapter [55$\ddagger$ FLUX 0.561 0.725 0.351
Diptych Prompting FLUX 0.689 0.758 0.344
@@ 4.3 Ablation Studies
We also assess the impact of varying the rescaling factor $\lambda$ on reference attention enhancement. Rescaling attention weights between the right panel query and the left panel key helps to capture fine details, thereby improving subject alignment metrics. However, using too high values introduces excessive inductive bias, causing abnormal attention weights that negatively impact performance. Qualitative transitions with respect to $\lambda$ can be verified in appendix.
@@ 4.4 Applications
Stylized Image Generation We extend our method beyond subject images and perform stylized image generation using various style images as references. Using style images and prompts in StyleDrop [43], we employ Diptych Prompting as same, but replace the subject name with the term ‘style’ in diptych text and without attention enhancement ($\lambda=1$) for referencing only the stylistic elements except the content. Diptych Prompting successfully generates the stylistic image reflecting the style of the reference as shown in Fig. 6, and the quantitative comparisons are available in appendix.
@@ Acknowledgement
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [No. 2022R1A3B1077720], Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [NO.RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), NO. RS-2022-II220959], the BK21 FOUR program of the Education and Research Program for Future ICT Pioneers, Seoul National University in 2024, Samsung Electronics (IO221213-04119-01), and a grant from the Yang Young Foundation.
Supplementary Material
A Baselines
@@
Figure S1: DreamBooth Comparisons. Quantitative comparisons to DreamBooth-LoRA with various rank values.
B Subject-Driven Text-to-Image Generation
@@ B.2 Comparison with Fine-Tuning-Based Method
To provide a more comprehensive comparison, we also compare with DreamBooth [41], a representative fine-tuning-based method. For efficient training, we attach a LoRA adapter to the pre-trained FLUX and perform fine-tuning by training only the LoRA adapter while freezing the FLUX. We train for 300 steps using the Adam optimizer with a learning rate of $1\times 10^{-4}$. Additionally, to compare different fine-tuning model capacities, we adjusted the rank of the LoRA adapter and conducted comparative experiments using the same metrics (DINO, CLIP-I, CLIP-T). The results are presented in Fig. S1, where our Diptych Prompting demonstrates superior performance across various model capacities.
@@ B.3 Additional Results
We include additional samples of Diptych Prompting in Fig. S2 and Fig. S3 for diverse objects and contexts. As demonstrated in the results, our methodology achieves high-quality image generation and satisfies both subject alignment and text alignment in a zero-shot manner by leveraging FLUX’s capabilities. Notably, this is accomplished without any specialized training for subject-driven text-to-image generation. We also note that the fine details in the target subject are well reflected in the generated results, even for challenging subjects that previous zero-shot methods struggled with (e.g., robot toy, ‘Bon appetit’ bowl).
C Human Preference Study
@@ D Diptych Generation
Table S1: Diptych Generation Comparisons. Quantitative comparisons of the diptych generation capabilities of various TTI models based on the total number of parameters, including the autoencoder, main network, and text encoder.
@@
Our framework relies on the emerging property of the large-scale TTI model, FLUX, particularly its strong understanding of diptych property and the ability to represent diptych accurately. We verify this by synthesizing a total of $2100$ diptychs, using $20$ objects, each with a pair of two random prompts for each panel among $15$ prompts, and comparing the diptych generation performance with those of other previous TTI models. The prompt for diptych generation follows the setup mentioned in Sec. 3.1 of the main paper. We assessed the quality of each diptych by evaluating the interrelation and text alignment of each panel. This is measured through splitting the generated image in half and measuring DINO and CLIP-I scores between each panel, as well as the CLIP-T score between each panel and its description. The results are shown in Tab. S1, in which the diptych generation performance and total number of parameters including the autoencoder, main network, and text encoders are reported. These results exhibit the superior diptych generation capability of FLUX, where smaller models are insufficient. This allows us to extend to inpainting and propose a zero-shot subject-driven text-to-image generation method via diptych inpainting-based interpretation.
E Background Removal Ablation
@@
We provide additional samples for the ablation study conducted with and without the background removal process $G_{\text{seg}}$ in Fig. S4. Consistent with the findings in the main paper, including the background leads to content leakage, where irrelevant elements such as background, pose, and location are mirrored in the generated results. This hinders the accurate reflection of the desired context described by the text and reduces diversity in pose and location. In contrast, removing the background and retaining only the subject information in the reference image on the left panel allows the generated outputs to better align with the desired context while exhibiting greater diversity in pose and location.”
F Reference Attention Enhancement Ablation
@@
We further present the actual sample quality variations according to the reference attention rescaling factor $\lambda$ values to support the quantitative ablations in the main paper. These variations are visualized in Fig. S5. As seen in the qualitative results, the absence of reference attention enhancement ($\lambda=1.0$) can lead to a loss of fine details of the subject, resulting in subtle discrepancies such as the left eye of the backpack dog, the patch on its right eye, the fur color on the dog’s face, or the texture of the bear plushie’s fur. As the $\lambda$ value increases, these missed details are better preserved, leading to improved subject alignment performance. However, excessive enhancement can negatively impact the quality of the generated images, causing the subject to appear slightly blurred or exhibit minor color shifts.
G Stylized Image Generation
@@
For stylized image generation, Diptych Prompting places the style image in the left panel and inpaints the right panel using the text prompt “A diptych with two side-by-side images of same style. On the left, {original image description}. On the right, replicate this style exactly but as {target image description} without attention enhancement ($\lambda=1.0$) for referencing only the stylistic elements except the content. Additional samples are provided in Fig. S6. Beyond the qualitative results, we also include quantitative comparisons using the same metrics (DINO, CLIP-T, CLIP-I) applied to a total of $2000$ generated images in Tab. S2. These images include $4$ samples per prompt and per style image, across $25$ prompts and $20$ style images collected from previous work [43]. As shown in the result, our method demonstrates comparable results to existing zero-shot style transfer methods specialized in stylized image generation, further proving the versatility of our approach.
@@
Table S2: Stylized Image Generation Comparisons. Quantitative comparisons of stylized image generation with previous zero-shot methods.
H Subject-Driven Image Editing
Diptych Prompting is extended to the subject-driven image editing by placing the reference subject image on the left panel and the editing target image on the right panel in the incomplete diptych. By masking only the desired area in the right panel and applying diptych inpainting, the reference subject from the left panel is generated in the masked region on the right panel, resulting in the subject-driven image editing. Following the previous work [54], we conduct the subject-driven image editing with selected images from a subset of the MSCOCO [26] validation dataset, in which each image contains a bounding box and the bounding box is smaller than half of image size. We applied masking to the inside of the bounding box, enabling the generation of the reference subject within the specified region. More samples of various subjects and editing target images are available in Fig. S7.
I Limitations
@@
Figure S2: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting.
Figure S3: Subject-Driven Text-to-Image Generation. More samples of subject-driven text-to-image generation using Diptych Prompting..
Figure S4: $\bm{G_{\text{seg}}}$ Ablation. Qualitative comparisons with and without the background removal process.
Figure S5: $\bm{\lambda}$ Ablation. Qualitative transitions according to the varying $\lambda$ values. we control the $\lambda$ from $1.0$ (without reference attention enhancement) to $1.5$. For a detailed view, please zoom in.
Figure S6: Stylized Image Generation. More samples of stylized image generation using Diptych Prompting.
Figure S7: Subject-Driven Image Editing. More samples of subject-driven image editing using Diptych Prompting.